OpenStudy (anonymous):

express the radical in simplified form negative square root of 300

OpenStudy (zzr0ck3r):

\(-\sqrt{300} = -\sqrt{3}\sqrt{100}=-10\sqrt{3}\)

OpenStudy (zzr0ck3r):

look for a nice square, then pull it out...

OpenStudy (anonymous):

thank you